OpenStudy (anonymous):

An artist has 30 kg of an alloy that is 45% copper. How many kilograms of a second alloy that is 70% copper should be mixed with the first alloy to get a new alloy that is 60% copper?

OpenStudy (anonymous):

Let x be the mass of second alloy, mass of copper in first alloy is 0.45*30 = 13.5 kg. (x*0.7)+13.5 = 0.6(30+x) , 0.1x = 18 - 13.5 = 4.5. => x = 45 kg. 45 kg of new alloy must be mixed.